# Distributed High-Dimensional Quantile Regression: Estimation Efficiency and Support Recovery

### v1

@@ A.2 Proof of Proposition 1 and Propositin 2
$\left\|\left(\widehat{\boldsymbol{D}}_{1,b,S^{c}\times S}-\boldsymbol{I}_{S^{c}\times S}\right)\left(\widehat{\boldsymbol{D}}_{1,b,S\times S}^{-1}-\boldsymbol{I}_{S\times S}^{-1}\right)\right\|_{\infty}=\mathcal{O}_{\mathbb{P}}\left(s^{3/2}\left(\frac{s\log n}{n}\right)^{2/3}\right)$ (38)
@@
$\left\|\boldsymbol{I}_{S^{c}\times S}\left(\widehat{\boldsymbol{D}}_{1,b,S\times S}^{-1}-\boldsymbol{I}_{S\times S}^{-1}\right)\right\|_{\infty}\leq s\left\|\boldsymbol{I}_{S^{c}\times S}\right\|_{op}\left\|\widehat{\boldsymbol{D}}_{1,b,S\times S}^{-1}-\boldsymbol{I}_{S\times S}^{-1}\right\|_{op}=\mathcal{O}_{\mathbb{P}}\left(s\left(\frac{s\log n}{n}\right)^{1/3}\right)$ (39)
@@
$\left\|\left(\widehat{\boldsymbol{D}}_{1,b,S^{c}\times S}-\boldsymbol{I}_{S^{c}\times S}\right)\boldsymbol{I}_{S\times S}^{-1}\right\|_{\infty}\leq s\left\|\left(\widehat{\boldsymbol{D}}_{1,b,S^{c}\times S}-\boldsymbol{I}_{S^{c}\times S}\right)\right\|_{op}\left\|\boldsymbol{I}^{-1}_{S\times S}\right\|_{op}=\mathcal{O}_{\mathbb{P}}\left(s\left(\frac{s\log n}{n}\right)^{1/3}\right)$ (40)
@@
For the second term, by the fact that $|\widetilde{Z}_{S}|_{\infty}\leq 1$, we have
$\left|\boldsymbol{M}_{2}\right|_{\infty}\leq\left\|\widehat{\boldsymbol{D}}_{1,b,S^{c}\times S}\widehat{\boldsymbol{D}}_{1,b,S\times S}^{-1}\right\|_{\infty}\left|\widetilde{Z}_{S}\right|_{\infty}\leq 1-\alpha/2.$ (42)
@@
This complete the proof. ∎
@@
where $C_{2}$ is a constant. Together with the assumption $s^{3}\log n/n=\mathcal{O}(1)$, we have
$\left\|\boldsymbol{I}_{S\times S}^{-1}\right\|_{\infty}\left|\boldsymbol{N}_{2}\right|_{\infty}\leq\frac{1}{2}\left|\widetilde{\boldsymbol{\beta}}_{S}-\boldsymbol{\beta}^{*}_{S}\right|_{\infty}$ (49)
with probablity tending to one. From the proof of Theorem 3.7, we can obtain that
@@
where $C_{3}$ is a constant, and $\eta=\max\left\{\left(\frac{s\log n}{n}\right)^{1/3},\left(\frac{s\log N}{n}\right)^{1/2}\right\}$. Furthermore, according to Proposition A.5, we have
@@ Appendix B Comparison to the Competitors
Space Aspects. The space complexity of DHSQR method is provided in Section 2.3 of the main text. Similarly, all other algorithms exhibit the same spatial complexity of $\mathcal{O}\left(np+p^{2}\right)$. While our algorithm demonstrates strong performance, it remains comparable to other methods in terms of space complexity. We prioritize efficient storage utilization, avoiding unnecessary memory overhead compared to competing algorithms.
@@
Table 1: Comparison of theoretical properties of different methods
statistical convergence Support recovery Heterogeneity
@@ C.1 Real Data Analysis
where $\widehat{y}_{i}$ represents the predictions made by each model. The results of the five different methods at quantile levels $\tau\in{0.5,0.75,0.9}$ are presented in Table 2. From the results, it can be observed that our DHSQR estimator exhibits smaller prediction errors compared to other distributed estimators and is very close to the performance of the single-machine estimator, Pooled DHSQR. This further underscores the method’s excellent performance on HIV data. It is noteworthy that the performance of the DREL estimator is even worse than the one-step Avg-DC estimator, indicating the presence of heteroscedasticity in the data. Clearly, the DREL algorithm is not suitable for analyzing this type of data.
@@ C.2.1 Sensitivity Analysis for the Bandwidth
Table 4: The $F_{1}$-score and $\ell_{2}$-error of the DHSQR under choices of bandwidth constant $c_{h},c_{b}$. (The standard deviation is given in parentheses)
@@ C.2.2 Computation Time Comparison
Table 5: the $F_{1}$ score, $\ell_{2}$ error, and Time computation of five estimators under different total sample size $N$. (The standard deviation is given in parentheses)
@@
As is shown in Table 5, We can see that our DHSQR has the fastest single iteration time, followed by DPQP, which maintains the same order of magnitude, then DREL and Pooled DHSQR, REL and Pooled DHSQR maintain the same order of magnitude, and Avg-DC is the slowest.
@@ C.2.3 Effect of Sample Size and Local Sample Size
Table 6: The $F_{1}$-score and $\ell_{2}$-error of the DHSQR, Pooled DHSQR, DPQR, DREL, and Avg-DC estimator (at iteration 10) under different sample size $N$ and noises are generated from normal distribution for the homoscedastic error case. The local sample size is fixed to $n=500$. (The standard deviation is given in parentheses)
@@
Table 7: The $F_{1}$-score and $\ell_{2}$-error of the DHSQR, Pooled DHSQR, DPQR, DREL, and Avg-DC estimator (at iteration 10) under different sample size $N$ and noises are generated from normal distribution for the Heteroscedastic error case. The local sample size is fixed to $n=500$. (The standard deviation is given in parentheses)
@@
Table 8: The $\ell_{2}$-error, precision, and recall of the DHSQR, Pooled DHSQR, DPQR, DREL, and Avg-DC estimator (at iteration 10) under different sample size $N$ and local sample size $n$. Noises are generated from a normal distribution for the Homoscedasticity error case. (The standard deviation is given in parentheses)
@@
Table 9: The $\ell_{2}$-error, precision, and recall of the DHSQR, Pooled DHSQR, DPQR, DREL, and Avg-DC estimator (at iteration 10) under different sample size $N$ and local sample size $n$. Noises are generated from a normal distribution for the Heteroscedasticity error case. (The standard deviation is given in parentheses)
@@
C.2.4 Additional Experiments for $\tau$
In this section, we provide some additional experiment results using quantile level $\tau=0.3$ and $\tau=0.7$. The results are reported in Tables. The observations are similar to the case of $\tau=0.5$ in Section 4.2. * denote Failure convergence.
Table 10: The $\ell_{2}$-error, precision, and recall of the DHSQR, pooled DHSQR, DPQR, DREL, and Avg-DC estimator (at iteration 10) under different sample size $N$ and local sample size $n$. Noises are generated from a normal distribution for the homoscedastic error case. (The standard deviation is given in parentheses, $\tau=0.3$)
@@ C.2.4 Additional Experiments for $\tau$
Table 11: The $\ell_{2}$-error, precision, and recall of the DHSQR, pooled DHSQR, DPQR, DREL, and Avg-DC estimator (at iteration 10) under different sample size $N$ and local sample size $n$. Noises are generated from a normal distribution for the heteroscedastic error case. (The standard deviation is given in parentheses, $\tau=0.3$)
@@
Table 12: The $\ell_{2}$-error, precision, and recall of the DHSQR, pooled DHSQR, DPQR, DREL, and Avg-DC estimator (at iteration 10) under different sample size $N$ and local sample size $n$. Noises are generated from a normal distribution for the homoscedastic error case. (The standard deviation is given in parentheses, $\tau=0.7$)
@@
Table 13: The $\ell_{2}$-error, precision, and recall of the DHSQR, pooled DHSQR, DPQR, DREL, and Avg-DC estimator (at iteration 10) under different sample size $N$ and local sample size $n$. Noises are generated from a normal distribution for the heteroscedastic error case. (The standard deviation is given in parentheses, $\tau=0.7$)
@@
C.2.5 Additional Experiments for $\boldsymbol{\beta}^{*}$ setting
In this section, we provide some additional experiment results using decaying sequence setting of nonzero parameters. In the case of Normal Heteroscedasticity, with $\tau=0.7,n=500,I=10$, and we set the true parameter as
@@ C.2.5 Additional Experiments for $\boldsymbol{\beta}^{*}$ setting
Other settings align with section 4. The average results from 100 replicates are summarized as follows Table.
Table 14: The $\ell_{2}$-error, precision, and recall of the DHSQR, pooled DHSQR, DPQR, DREL, and Avg-DC estimator (at iteration 10) under different sample size $N$ and local sample size $n$. Noises are generated from a normal distribution for the heteroscedastic error case. (The standard deviation is given in parentheses, $\tau=0.7$)

### v3

@@ A.1 Some Technical Lemmas
$\left\|\left(\widehat{\boldsymbol{D}}_{1,b,S^{c}\times S}-\boldsymbol{I}_{S^{c}\times S}\right)\left(\widehat{\boldsymbol{D}}_{1,b,S\times S}^{-1}-\boldsymbol{I}_{S\times S}^{-1}\right)\right\|_{\infty}=\mathcal{O}_{\mathbb{P}}\left(s^{3/2}\left(\frac{s\log n}{n}\right)^{2/3}\right).$ (39)
@@
$\left\|\boldsymbol{I}_{S^{c}\times S}\left(\widehat{\boldsymbol{D}}_{1,b,S\times S}^{-1}-\boldsymbol{I}_{S\times S}^{-1}\right)\right\|_{\infty}\leq s\left\|\boldsymbol{I}_{S^{c}\times S}\right\|_{op}\left\|\widehat{\boldsymbol{D}}_{1,b,S\times S}^{-1}-\boldsymbol{I}_{S\times S}^{-1}\right\|_{op}=\mathcal{O}_{\mathbb{P}}\left(s\left(\frac{s\log n}{n}\right)^{1/3}\right),$ (40)
@@
$\left\|\left(\widehat{\boldsymbol{D}}_{1,b,S^{c}\times S}-\boldsymbol{I}_{S^{c}\times S}\right)\boldsymbol{I}_{S\times S}^{-1}\right\|_{\infty}\leq s\left\|\left(\widehat{\boldsymbol{D}}_{1,b,S^{c}\times S}-\boldsymbol{I}_{S^{c}\times S}\right)\right\|_{op}\left\|\boldsymbol{I}^{-1}_{S\times S}\right\|_{op}=\mathcal{O}_{\mathbb{P}}\left(s\left(\frac{s\log n}{n}\right)^{1/3}\right).$ (41)
@@
For the second term, by the fact that $|\widetilde{\boldsymbol{Z}}_{S}|_{\infty}\leq 1$, we have
$\left|\boldsymbol{M}_{2}\right|_{\infty}\leq\left\|\widehat{\boldsymbol{D}}_{1,b,S^{c}\times S}\widehat{\boldsymbol{D}}_{1,b,S\times S}^{-1}\right\|_{\infty}\left|\widetilde{\boldsymbol{Z}}_{S}\right|_{\infty}\leq 1-\alpha/2.$ (43)
@@
This completes the proof. ∎
@@
where $C_{2}$ is a constant. Together with the assumption $s^{3}\log n/n=o(1)$, we have
$\left\|\boldsymbol{I}_{S\times S}^{-1}\right\|_{\infty}\left|\boldsymbol{N}_{2}\right|_{\infty}\leq\frac{1}{2}\left|\widetilde{\boldsymbol{\beta}}_{S}-\boldsymbol{\beta}^{*}_{S}\right|_{\infty},$ (50)
with probability tending to one. From the proof of Theorem 3.7, we can obtain that
@@
where $C_{3}$ is a constant, and $\eta=\max\left\{\left(\frac{s\log n}{n}\right)^{1/3},\left(\frac{s\log N}{n}\right)^{1/2}\right\}$.
Furthermore, according to Proposition A.5, we have
@@ Appendix B Comparison to the Competitors
Space Aspects. The space complexity of the DHSQR method is provided in Section 2.3 of the main text. Similarly, all other algorithms exhibit the same spatial complexity of $\mathcal{O}\left(np+p^{2}\right)$. While our algorithm demonstrates strong performance, it remains comparable to other methods in terms of space complexity. We prioritize efficient storage utilization, avoiding unnecessary memory overhead compared to competing algorithms.
@@
Table 1: Comparison of theoretical properties of different methods.
Method Statistical convergence Support recovery Heterogeneity
@@ C.1 Real Data Analysis
where $\widehat{y}_{i}$ represents the predictions made by each model. The results of the five different methods at quantile levels $\tau\in\{0.5,0.75,0.9\}$ are presented in Table 2. From the results, it can be observed that our DHSQR estimator exhibits smaller prediction errors compared to other distributed estimators and is very close to the performance of the single machine estimator, Pooled DHSQR. This further underscores the method’s excellent performance on HIV data. It is noteworthy that the performance of the DREL estimator is even worse than the one-step Avg-DC estimator, indicating the presence of heteroscedasticity in the data. Clearly, the DREL algorithm is not suitable for analyzing this type of data.
@@ C.2.1 Sensitivity Analysis for the Bandwidth
Table 4: The $F_{1}$-score and $\ell_{2}$-error of the DHSQR under different choices of bandwidth constant $c_{h},c_{b}$. (The standard deviation is given in parentheses.)
@@ C.2.2 Computation Time Comparison
Table 5: the $F_{1}$ score, $\ell_{2}$ error, and Time computation of five estimators under different total sample sizes $N$. (The standard deviation is given in parentheses.)
@@
As is shown in Table 5, We can see that our DHSQR has the fastest single iteration time, followed by DPQP, which maintains the same order of magnitude, then DREL and Pooled DHSQR, REL, and Pooled DHSQR maintain the same order of magnitude, and Avg-DC is the slowest.
@@ C.2.3 Effect of Sample Size and Local Sample Size
Table 6: The $F_{1}$-score and $\ell_{2}$-error of the DHSQR, Pooled DHSQR, DPQR, DREL, and Avg-DC estimator (at iteration 10) under different sample size $N$. Noises are generated from a normal distribution for the homoscedastic error case. The local sample size is fixed to $n=500$. The quantile level is fixed $\tau=0.5$ and the iteration is fixed $T=10$. (The standard deviation is given in parentheses.)
@@
Table 7: The $F_{1}$-score and $\ell_{2}$-error of the DHSQR, Pooled DHSQR, DPQR, DREL, and Avg-DC estimator (at iteration 10) under different sample size $N$. Noises are generated from a normal distribution for the heteroscedastic error case. The local sample size is fixed to $n=500$. The quantile level is fixed $\tau=0.5$ and the iteration is fixed $T=10$. (The standard deviation is given in parentheses.)
@@
Table 8: The $\ell_{2}$-error, precision, and recall of the DHSQR, Pooled DHSQR, DPQR, DREL, and Avg-DC estimator under different sample size $N$ and local sample size $n$. Noises are generated from a normal distribution for the homoscedastic error case. The quantile level is fixed $\tau=0.5$ and the iteration is fixed $T=10$. (The standard deviation is given in parentheses.)
@@
Table 9: The $\ell_{2}$-error, precision, and recall of the DHSQR, Pooled DHSQR, DPQR, DREL, and Avg-DC estimator under different sample size $N$ and local sample size $n$. Noises are generated from a normal distribution for the heteroscedastic error case. The quantile level is fixed $\tau=0.5$ and the iteration is fixed $T=10$. (The standard deviation is given in parentheses.)
@@
C.2.4 Additional Experiments for $\tau=\{0.3,0.7\}$
In this section, we provide some additional experiment results using quantile level $\tau=0.3$ and $\tau=0.7$. The results are reported in tables. The observations are similar to the case of $\tau=0.5$ in Section 4.2. * indicates failure to converge.
Table 10: The $\ell_{2}$-error, precision, and recall of the DHSQR, Pooled DHSQR, DPQR, DREL, and Avg-DC estimator under different sample size $N$ and local sample size $n$. Noises are generated from a normal distribution for the homoscedastic error case. The quantile level is fixed $\tau=0.3$ and the iteration is fixed $T=10$. (The standard deviation is given in parentheses.)
@@
Table 11: The $\ell_{2}$-error, precision, and recall of the DHSQR, Pooled DHSQR, DPQR, DREL, and Avg-DC estimator under different sample size $N$ and local sample size $n$. Noises are generated from a normal distribution for the heteroscedastic error case. The quantile level is fixed $\tau=0.3$ and the iteration is fixed $T=10$. (The standard deviation is given in parentheses.)
@@
Table 12: The $\ell_{2}$-error, precision, and recall of the DHSQR, Pooled DHSQR, DPQR, DREL, and Avg-DC estimator under different sample size $N$ and local sample size $n$. Noises are generated from a normal distribution for the homoscedastic error case. The quantile level is fixed $\tau=0.7$ and the iteration is fixed $T=10$. (The standard deviation is given in parentheses.)
@@
Table 13: The $\ell_{2}$-error, precision, and recall of the DHSQR, Pooled DHSQR, DPQR, DREL, and Avg-DC estimator under different sample size $N$ and local sample size $n$. Noises are generated from a normal distribution for the heteroscedastic error case. The quantile level is fixed $\tau=0.7$ and the iteration is fixed $T=10$. (The standard deviation is given in parentheses.)
@@
C.2.5 Additional Experiments for the Decaying Sequence Setting of Nonzero Parameters
In this section, we provide some additional experiment results using the decaying sequence setting of nonzero parameters. We consider the heteroscedastic error case with normal distribution, and $\tau=0.7,n=500,T=10$. Here, we set the true parameter as
@@ C.2.5 Additional Experiments for the Decaying Sequence Setting of Nonzero Parameters
Other settings align with those in Section 4. The average results from 100 replicates are summarized in Table 14.
Table 14: The $\ell_{2}$-error, precision, and recall of the DHSQR, pooled DHSQR, DPQR, DREL, and Avg-DC estimator under different sample size $N$ and local sample size $n$. Noises are generated from a normal distribution for the heteroscedastic error case. The quantile level is fixed $\tau=0.7$ and the iteration is fixed $T=10$. (The standard deviation is given in parentheses.)